# On the solution of the Collatz problem

### v26

@@ 3.2. Properties of Collatz’s Sequences
where $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m+1$, let $n_{m+1}(m,t)=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})$ where $r_{1}\geq 1$, $0\leq i_{1}\leq m_{1}$, and $n_{i_{1}}(m_{1},t_{1})$ is odd and not divisible by $3$.
Then for $i=0,1,2,\ldots,m$ we have
@@
$=(2/3)(1-\frac{1}{2n_{i+1}(m,t)})n_{i+1}(m,t)$
$=(2/3)^{2}\prod_{k=i}^{i+1}(1-\frac{1}{2n_{k+1}(m,t)})n_{i+2}(m,t)=\cdots$
$=(2/3)^{m+1-i}\prod_{k=i}^{m}(1-\frac{1}{2n_{k+1}(m,t)})n_{m+1}(m,t)$
$=(2/3)^{m+1-i}2^{r_{1}}\prod_{k=i}^{m}(1-\frac{1}{2n_{k+1}(m,t)})n_{i_{1}}(m_{1},t_{1}).$
@@
Lemma 3.7 (On cycle in a Collatz’s Sequence).
No Collatz cycle contains any odd number greater than 1 in a Collatz’s Sequence $S_{c}(m,t)$.
@@
Second, if $i=m=0$ and $t>0$ or $m-i=0$ and $m>0$ then there are
@@
If $i>0$ then there are $13^{i}t>$ 1 and $(2\times 2^{m+1-i}-3^{m+1-i})3^{i}t\neq 1$, so that $r_{c}(m,t,i,1)\neq 1$, else for $i=0$ only if $m=0$ and $t=1$ then there is $(2\times 2^{m+1}-3^{m+1})t=1$, i.e., only if $n_{i}(m,t)=n_{0}(0,1)=1$ and $r_{1}=1$ then there is $r_{c}(m,t,i,1)=1$.
Hence no Collatz cycle contains any odd number greater than 1 in $S_{c}(m,t)$.
@@
$=\prod_{j=0}^{1}(2/3)^{m_{j}+1-i_{j}}2^{r_{j+1}}\prod_{k=i_{j}}^{m_{j}}(1-\frac{1}{2n_{k+1}(m_{j},t_{j})})n_{i_{2}}(m_{2},t_{2})$
$=\prod_{j=0}^{2}(2/3)^{m_{j}+1-i_{j}}2^{r_{j+1}}\prod_{k=i_{j}}^{m_{j}}(1-\frac{1}{2n_{k+1}(m_{j},t_{j})})n_{i_{3}}(m_{3},t_{3})$
$=\cdots=\prod_{j=0}^{s-1}(2/3)^{m_{j}+1-i_{j}}2^{r_{j+1}}\prod_{k=i_{j}}^{m_{j}}(1-\frac{1}{2n_{k+1}(m_{j},t_{j})})n_{i_{s}}(m_{s},t_{s}).$
@@
Lemma 3.9 (On cycle in a series of Collatz’s Sequences).
No Collatz cycle contains any odd number greater than 1 in a series of Collatz’s Sequences $S_{c}(s)$.
@@
If $s=1$ then by Lemma 3.7 no Collatz cycle contains any odd number greater than 1 in a Collatz’s Sequence $S_{c}(m,t)$.
@@
let an odd number $n_{i}(m,t)$ greater than 1 be in the first Collatz’s Sequence $S_{c}(m,t)$ where $0\leq i\leq m$,
$\sigma(\zeta)=\sum_{j=1}^{\zeta}r_{j}\texttt{, }\sigma(k,\zeta)=\sum_{j=k}^{\zeta}r_{j}\texttt{, and }\eta_{\xi}=\sum_{j=0}^{\xi}(m_{j}-i_{j}+1)$
for $\xi=0,1,2,\cdots,s-1$. Then by Lemma 3.8, with $0\leq i\leq m$ and $i_{0}=i$ we have
@@
Thus if $r_{c}(m,t,i,s)=1$ then there must be $n_{i_{s}}(m_{s},t_{s})=n_{i}(m,t)$ and
$\prod_{j=0}^{s-1}3^{m_{j}+1-i_{j}}<\prod_{j=0}^{s-1}2^{m_{j}+1-i_{j}}2^{r_{j+1}}\texttt{ or }3^{\eta}<2^{\eta+\sigma(s)}\texttt{ where }\eta=\eta_{s-1},$
and there must be at least one Collatz’s Sequence $S_{c}(m_{h},t_{h})\in S_{c}(s)$ where $0\leq h<s$, in which the element $n_{i_{h}}(m_{h},t_{h})$ satisfies $m_{h}-i_{h}+1\geq 2$, otherwise there shall be $n_{i_{s}}(m_{s},t_{s})<n_{i}(m,t)$ such that $r_{c}(m,t,i,s)>1$.
By the assumption $r_{c}(m,t,i,s)=1$ for $1<s<\infty$ and by Definition 1.1 there are $\eta$ linear equations
@@
where $n_{m_{j}+1}(m_{j},t_{j})=2^{r_{j+1}}n_{i_{j+1}}(m_{j+1},t_{j+1})$ for $j=0,1,2,\ldots,s-1$ and $n_{i_{s}}(m_{s},t_{s})=n_{i}(m,t)$.
Let rewrite these linear equations in the form of a matrix equation. Then there is $\mathbf{A}\mathbf{X}=\mathbf{B}$ where
$\mathbf{A}=\left(\begin{array}[]{ccccccccccccc}3&-2&0&\cdots&0&0&0&\cdots&0&0&0&\cdots&0\\ 0&3&-2&\cdots&0&0&0&\cdots&0&0&0&\cdots&0\\ 0&0&3&\cdots&0&0&0&\cdots&0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ 0&0&0&\cdots&3&-2^{1+r_{1}}&0&\cdots&0&0&0&\cdots&0\\ 0&0&0&\cdots&0&3&-2&\cdots&0&0&0&\cdots&0\\ 0&0&0&\cdots&0&0&3&\cdots&0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ 0&0&0&\cdots&0&0&0&\cdots&3&-2^{1+r_{2}}&0&\cdots&0\\ 0&0&0&\cdots&0&0&0&\cdots&0&3&-2&\cdots&0\\ 0&0&0&\cdots&0&0&0&\cdots&0&0&3&\cdots&0\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ -2^{1+r_{s}}&0&0&\cdots&0&0&0&\cdots&0&0&0&\cdots&3\\ \end{array}\right),$
$\mathbf{B}=-\left(\begin{array}[]{ccccccccccccc}1&1&\cdots&1&1&1&\cdots&1&\cdots&1&1&\cdots&1\\ \end{array}\right)^{T},$
and
$\mathbf{X}=\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{\eta}\\ \end{array}\right)=\left(\begin{array}[]{c}n_{i}(m,t)\\ n_{i+1}(m,t)\\ \vdots\\ n_{m}(m,t)\\ n_{i_{1}}(m_{1},t_{1})\\ n_{i_{1}+1}(m_{1},t_{1})\\ \vdots\\ n_{m_{1}}(m_{1},t_{1})\\ \vdots\\ n_{i_{s-1}}(m_{s-1},t_{s-1})\\ n_{i_{s-1}+1}(m_{s-1},t_{s-1})\\ \vdots\\ n_{m_{s-1}}(m_{s-1},t_{s-1})\\ \end{array}\right),$
i.e., $\mathbf{A}$ expressed as above or in Lemma 3.10 is a matrix with two dimensions $\eta\times\eta$, and $\mathbf{B}$ and $\mathbf{X}$ are vectors with a dimension $\eta$.
@@
(3.3) $Det\mathbf{A}=3^{\eta}-2^{\eta+\sigma(s)}<0\texttt{ and }-Det\mathbf{A}=2^{\eta+\sigma(s)}-3^{\eta}\geq 2^{3+\sigma(2)}-3^{3}\geq 5$
where it is obvious that $-Det\mathbf{A}$ can not be dividable by $3$ so that $-Det\mathbf{A}\neq 3$, and when $\eta>1$ or $\eta+\sigma(s)>2$ if $-Det\mathbf{A}=1$, i.e., if $4(2^{\eta+\sigma(s)-2}-1)=3(3^{\eta-1}-1)$, then first for $\eta=2k+2$ where $k\geq 0$, there should be $2(2^{\eta+\sigma(s)-2}-1)=3\sum_{i=0}^{2k}3^{i}$, but this equality can not hold since its left-hand value is even and its right-hand value is odd, so that $-Det\mathbf{A}\neq 1$, and next for $\eta=2k+3$ where $k\geq 0$, there should be $2^{\eta+\sigma(s)-2}-1=6\sum_{i=0}^{k}3^{2i}$, but this equality can not hold since its left-hand value is odd and its right-hand value is even, so that $-Det\mathbf{A}\neq 1$.
Then the matrix $\mathbf{A}$ is nonsingular and there is $\mathbf{A}^{-1}=\mathbf{A}^{*}/Det\mathbf{A}$ where $\mathbf{A}^{-1}$ and $\mathbf{A}^{*}$ are the inverse matrix and the adjoint matrix of the matrix $\mathbf{A}$, respectively, and
(3.4) $\mathbf{A}^{*}=\left(\begin{array}[]{ccccccc}A_{1,1}&A_{2,1}&A_{3,1}&\cdots&A_{\eta-2,1}&A_{\eta-1,1}&A_{\eta,1}\\ A_{1,2}&A_{2,2}&A_{3,2}&\cdots&A_{\eta-2,2}&A_{\eta-1,2}&A_{\eta,2}\\ A_{1,3}&A_{2,3}&A_{3,3}&\cdots&A_{\eta-2,3}&A_{\eta-1,3}&A_{\eta,3}\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ A_{1,\eta-2}&A_{2,\eta-2}&A_{3,\eta-2}&\cdots&A_{\eta-2,\eta-2}&A_{\eta-1,\eta-2}&A_{\eta,\eta-2}\\ A_{1,\eta-1}&A_{2,\eta-1}&A_{3,\eta-1}&\cdots&A_{\eta-2,\eta-1}&A_{\eta-1,\eta-1}&A_{\eta,\eta-1}\\ A_{1,\eta}&A_{2,\eta}&A_{3,\eta}&\cdots&A_{\eta-2,\eta}&A_{\eta-1,\eta}&A_{\eta,\eta}\\ \end{array}\right).$
@@
and there must be different positive integers
$x_{k}=-\sum_{j=1}^{\eta}A_{j,k}/Det\mathbf{A}\texttt{ for }k=1,2,\ldots,\eta.$
By Lemma 3.10 the matrix $\mathbf{A}^{*}$ is determined. If each $x_{k}$ for $k=1,2,\ldots,\eta$ is a positive integer then each number $-\sum_{j=1}^{\eta}A_{j,k}/Det\mathbf{A}$ for $k=1,2,\ldots,\eta$ must be an exact division or each number $\sum_{j=1}^{\eta}A_{j,k}$ for $k=1,2,\ldots,\eta$ must be exactly dividable by $-Det\mathbf{A}$.
In this case, we can take any pair of $\sum_{j=1}^{\eta}A_{j,k}$ and $\sum_{j=1}^{\eta}A_{j,k+1}$ which are consecutive and must be exactly dividable by $-Det\mathbf{A}$ so that their difference must be also exactly dividable by $-Det\mathbf{A}$.
To make easy and clear, we take $\sum_{j=1}^{\eta}A_{j,1}$ and $\sum_{j=1}^{\eta}A_{j,2}$ which are consecutive. Let $\Delta_{1,2}$ denote their difference. Then by Formula 3.5, there are
$\Delta_{1,2}=\sum_{j=1}^{\eta}A_{j,2}-\sum_{j=1}^{\eta}A_{j,1}=A_{1,2}-A_{\eta,1}$
$=2^{\eta-1+\sigma(s)}-2^{\eta-1+\sigma(s-1)}=(2^{r_{s}}-1)2^{\eta-1+\sigma(s-1)}$
and the number $12^{r_{s}}-$ 1 must be exactly dividable by $-Det\mathbf{A}$. So there must be $r_{s}\geq 3$ since by Formula 3.3 there is $-Det\mathbf{A}\geq 5$.
By the assumption $r_{c}(m,t,i,s)=1$ for $1<s<\infty$, all positive odd integers $x_{k}$ for $k=1,2,\ldots,\eta$ form a Collatz cycle. Then we can take each positive odd integer $n_{i_{\xi}}(m_{\xi},t_{\xi})$ for $\xi=0,1,2,\ldots,s-1$ as the first element $x_{1}$ of the cycle, so that there must be $r_{j+1}\geq 3$ for $j=0,1,2,\ldots,s-1$. But it conflicts with the fact that by Lemma 3.5 a half of positive integers $r_{j+1}$ satisfy $r_{j+1}=1$ and a quarter of positive integers $r_{j+1}$ satisfy $r_{j+1}=2$ where $j=0,1,2,\ldots,s-1$. Thus the assumption $r_{c}(m,t,i,s)=1$ does not hold for each finite positive integer $s$ greater than 1 and less than $\infty$.
Hence no Collatz cycle contains any odd number greater than 1 in $S_{c}(s)$.
@@
Let $Det\mathbf{A}$ denote the determinant of a matrix $\mathbf{A}$, $\sigma(\zeta)=\sum_{j=1}^{\zeta}r_{j}$, $\sigma(k,\zeta)=\sum_{j=k}^{\zeta}r_{j}$, and $\eta_{\xi}=\sum_{j=0}^{\xi}(m_{j}-i_{j}+1)$ for $\xi=0,1,2,\cdots,s-1$ where $0\leq i_{j}\leq m_{j}$ and $\eta=\eta_{s-1}$. Then given a $\eta\times\eta$ nonsingular matrix
where $Det\mathbf{A}=3^{\eta}-2^{\eta+\sigma(s)}$, the adjoint matrix of the matrix $\mathbf{A}$ is
(3.5) $\mathbf{A}^{*}=\left(\begin{array}[]{ccccccc}A_{1,1}&A_{2,1}&A_{3,1}&\cdots&A_{\eta-2,1}&A_{\eta-1,1}&A_{\eta,1}\\ A_{1,2}&A_{1,1}&A_{2,1}&\cdots&A_{\eta-3,1}&A_{\eta-2,1}&A_{\eta-1,1}\\ A_{1,3}&A_{1,2}&A_{1,1}&\cdots&A_{\eta-4,1}&A_{\eta-3,1}&A_{\eta-2,1}\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ A_{1,\eta-2}&A_{1,\eta-3}&A_{1,\eta-4}&\cdots&A_{1,1}&A_{2,1}&A_{3,1}\\ A_{1,\eta-1}&A_{1,\eta-2}&A_{1,\eta-3}&\cdots&A_{1,2}&A_{1,1}&A_{2,1}\\ A_{1,\eta}&A_{1,\eta-1}&A_{1,\eta-2}&\cdots&A_{1,3}&A_{1,2}&A_{1,1}\\ \end{array}\right)$
where on the first collum of the matrix $\mathbf{A}^{*}$ there are
(3.6) $A_{1,k}=2^{\eta-k+1+\sigma(s)}3^{k-2}\texttt{ for }k=2,3,\ldots,\eta_{0}$
and for $\xi=0,1,2,\ldots,s-2$
(3.7) $A_{1,k}=2^{\eta-k+1+\sigma(\xi+2,s)}3^{k-2}\texttt{ for }k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1},$
and on the last collum of the matrix $\mathbf{A}^{*}$ there are
(3.8) $A_{\eta,k}=2^{\eta-k+\sigma(s-1)}3^{k-1}\texttt{ for }k=1,2,\ldots,\eta_{0}$
and for $\xi=0,1,2,\ldots,s-2$
(3.9) $A_{\eta,k}=2^{\eta-k+\sigma(\xi+2,s-1)}3^{k-1}\texttt{ for }k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1},$
and on the first row of the matrix $\mathbf{A}^{*}$ there are
(3.10) $A_{k,1}=2A_{k-1,1}/3=2^{k-1}3^{\eta-k}\texttt{ for }k=2,3,\ldots,\eta_{0}$
and for $\xi=0,1,2,\ldots,s-2$
(3.11) $A_{k,1}=2^{k-1+\sigma(\xi+1)}3^{\eta-k}\texttt{ for }k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\eta_{\xi+1},$
and on the $\kappa^{th}$ row of the matrix $\mathbf{A}^{*}$ where $\kappa=3,4,\ldots,\eta$ and positive integers $A_{k,\kappa}$ for $k=2,3,\ldots,\kappa-1$ there are
(3.12) $A_{k,\kappa}=2A_{k-1,\kappa}/3=2^{\eta-\kappa+k+\sigma(s)}3^{\kappa-k-1}\texttt{ for }k=2,3,\ldots,\min(\eta_{0},\kappa-1)$
and for $k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\kappa-1)$
(3.13) $A_{k,\kappa}=2^{k-1+\sigma(\xi+1)}3^{\eta-k}\texttt{ for }k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\kappa-1),$
and on the $\kappa^{th}$ row of the matrix $\mathbf{A}^{*}$ where $\kappa=2,3,\ldots,\eta-1$ and positive integers $A_{\kappa+k,\kappa}$ for $k=1,2,\ldots,\eta-\kappa$ there are
(3.14) $A_{\kappa+k,\kappa}=2A_{\kappa+k-1,\kappa}/3=2^{k}3^{\eta-1-k}\texttt{ for }k=1,2,\ldots,\min(\eta_{0},\eta-\kappa)$
and for $k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta_{\xi}+\eta-\kappa)$
(3.15) $A_{\kappa+k,\kappa}=2^{k+\sigma(\xi+1)}3^{\eta-1-k}\texttt{ for }k=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta_{\xi}+\eta-\kappa).$

### v32

@@ 3.2. Properties of Collatz’s Sequences
where $n_{i}(m,t)=2^{m+1-i}\times 3^{i}t-1$ for $i=0,1,2,\ldots,m+1$, let
$n_{m+1}(m,t)=2^{r_{1}}n_{i_{1}}(m_{1},t_{1})\texttt{ where }r_{1}\geq 1\texttt{, }0\leq i_{1}\leq m_{1},$
and $n_{i_{1}}(m_{1},t_{1})$ is odd and not divisible by $3$. Then for $i=0,1,2,\ldots,m$ we have
@@
$(2/3)(1-\frac{1}{2n_{i+1}(m,t)})n_{i+1}(m,t)$
$(2/3)^{2}\prod_{k=i}^{i+1}(1-\frac{1}{2n_{k+1}(m,t)})n_{i+2}(m,t)=\cdots$
$(2/3)^{m+1-i}\prod_{k=i}^{m}(1-\frac{1}{2n_{k+1}(m,t)})n_{m+1}(m,t)$
$(2/3)^{m+1-i}2^{r_{1}}\prod_{k=i}^{m}(1-\frac{1}{2n_{k+1}(m,t)})n_{i_{1}}(m_{1},t_{1}).$
@@
Lemma 3.7.
No nontrivial Collatz cycle exists in any Collatz’s Sequence $S_{c}(m,t)$.
@@
Second, if $i=m=0$ and $t>0$ or mi=0m-i=0 and $m>0$ then there are
@@
If $i>0$ then there are $3^{i}t>1$ and $(2\times 2^{m+1-i}-3^{m+1-i})3^{i}t\neq 1$, so that $r_{c}(m,t,i,1)\neq 1$, else for $i=0$ only if $m=0$ and $t=1$ then there is $(2\times 2^{m+1}-3^{m+1})t=1$, i.e., only if $n_{i}(m,t)=n_{0}(0,1)=1$ and $r_{1}=1$ then there is $r_{c}(m,t,i,1)=1$.
Hence no nontrivial Collatz cycle exists in $S_{c}(m,t)$.
@@
$\prod_{j=0}^{1}(2/3)^{m_{j}+1-i_{j}}2^{r_{j+1}}\prod_{k=i_{j}}^{m_{j}}(1-\frac{1}{2n_{k+1}(m_{j},t_{j})})n_{i_{2}}(m_{2},t_{2})$
$\prod_{j=0}^{2}(2/3)^{m_{j}+1-i_{j}}2^{r_{j+1}}\prod_{k=i_{j}}^{m_{j}}(1-\frac{1}{2n_{k+1}(m_{j},t_{j})})n_{i_{3}}(m_{3},t_{3})$
$\cdots=\prod_{j=0}^{s-1}(2/3)^{m_{j}+1-i_{j}}2^{r_{j+1}}\prod_{k=i_{j}}^{m_{j}}(1-\frac{1}{2n_{k+1}(m_{j},t_{j})})n_{i_{s}}(m_{s},t_{s}).$
@@
Lemma 3.9.
No nontrivial Collatz cycle exists in any series of Collatz’s Sequences $S_{c}(s)$ where $0<s<\infty$.
@@
If $s=1$ then by Lemma 3.7 no nontrivial Collatz cycle exists in any Collatz’s Sequence $S_{c}(m,t)$.
@@
Let an odd number $n_{i}(m,t)$ greater than 1 be in the first Collatz’s Sequence $S_{c}(m,t)$ where $0\leq i\leq m$, and by Definition 1.8 there are $\eta=\eta_{s}$, $\eta_{0}=0$,
$\sigma(\zeta)=\sum_{j=1}^{\zeta}r_{j}\texttt{, }\sigma(\xi,\zeta)=\sum_{j=\xi}^{\zeta}r_{j}\texttt{, and }\eta_{\xi}=\sum_{j=0}^{\xi-1}(m_{j}-i_{j}+1)\texttt{ where }0\leq i_{j}\leq m_{j}$
for $\xi=1,2,\cdots,s$. Then by Lemma 3.8, with $0\leq i\leq m$ and $i_{0}=i$ we have
@@
Thus if $r_{c}(m,t,i,s)\neq 1$ then no nontrivial Collatz cycle exists in $S_{c}(s)$ else if $r_{c}(m,t,i,s)=1$ then there must be $n_{i_{s}}(m_{s},t_{s})=n_{i}(m,t)$,
$\prod_{j=0}^{s-1}3^{m_{j}+1-i_{j}}<\prod_{j=0}^{s-1}2^{m_{j}+1-i_{j}}2^{r_{j+1}}\texttt{ or }3^{\eta}<2^{\eta+\sigma(s)}\texttt{ where }\eta=\eta_{s},$
and there must exist at least one Collatz’s Sequence $S_{c}(m_{h},t_{h})\in S_{c}(s)$ where $0\leq h<s$, in which the element $n_{i_{h}}(m_{h},t_{h})$ satisfies $m_{h}-i_{h}+1\geq 2$, i.e., there must be $\eta>s$, otherwise there shall be $n_{i_{s}}(m_{s},t_{s})<n_{i}(m,t)$ such that $r_{c}(m,t,i,s)>1$.
By the assumption $r_{c}(m,t,i,s)=1$ for $1<s<\infty$ and by Definition 1.1 for $j=0,1,2,\ldots,s-1$ there are $\eta$ linear equations
@@
where $n_{m_{j}+1}(m_{j},t_{j})=2^{r_{j+1}}n_{i_{j+1}}(m_{j+1},t_{j+1})$ and $n_{i_{s}}(m_{s},t_{s})=n_{i}(m,t)$.
Let rewrite these linear equations in the form of a matrix equation. Then there is $\mathbf{A}\mathbf{X}=\mathbf{B}$ where $\mathbf{A}$ is defined by Definition 1.9 with two dimensions $\eta\times\eta$,
$\mathbf{B}=-\left(\begin{array}[]{cccc}1&1&\cdots&1\\ \end{array}\right)^{T}\texttt{ and }\mathbf{X}=\left(\begin{array}[]{cccc}x_{1}&x_{2}&\cdots&x_{\eta}\\ \end{array}\right)^{T}$
are vectors with a dimension $\eta$ where for $j=0,1,2,\ldots,s-1$
$x_{\eta_{j}+1}=n_{i_{j}}(m_{j},t_{j})\texttt{, }x_{\eta_{j}+2}=n_{i_{j}+1}(m_{j},t_{j})\texttt{, }\ldots\texttt{ }x_{\eta_{j+1}}=n_{m_{j}}(m_{j},t_{j}).$
@@
(3.3) $Det\mathbf{A}=3^{\eta}-2^{\eta+\sigma(s)}<0\texttt{, }-Det\mathbf{A}=2^{\eta+\sigma(s)}-3^{\eta}\geq 5\texttt{, and }\eta>s$
where it is obvious that $Det\mathbf{A}$ can not be dividable by $3$ so that $Det\mathbf{A}\neq 3$, and when $\eta>1$ or $\eta+\sigma(s)>2$ if $Det\mathbf{A}=1$, i.e., if $4(2^{\eta+\sigma(s)-2}-1)=3(3^{\eta-1}-1)$, then first for $\eta=2k+2$ where $k\geq 0$, there should be $2(2^{\eta+\sigma(s)-2}-1)=3\sum_{i=0}^{2k}3^{i}$, but this equality can not hold since its left-hand value is even and its right-hand value is odd, so that $Det\mathbf{A}\neq 1$, and next for $\eta=2k+3$ where $k\geq 0$, there should be $2^{\eta+\sigma(s)-2}-1=6\sum_{i=0}^{k}3^{2i}$, but this equality can not hold since its left-hand value is odd and its right-hand value is even, so that $Det\mathbf{A}\neq 1$.
Then the matrix $\mathbf{A}$ is nonsingular and there is $\mathbf{A}^{-1}=\mathbf{A}^{*}/Det\mathbf{A}$ where $\mathbf{A}^{-1}$ and $\mathbf{A}^{*}$ are the inverse matrix and the adjoint matrix of the matrix $\mathbf{A}$, respectively.
@@
and for $\kappa=1,2,\ldots,\eta$ there must be different positive odd integers
(3.4) $x_{\kappa}=-\sum_{j=1}^{\eta}A_{j,\kappa}/Det\mathbf{A}=-A_{\kappa}/Det\mathbf{A}\texttt{ where }A_{\kappa}=\sum_{j=1}^{\eta}A_{j,\kappa}.$
Thus for $\kappa=1,2,\ldots,\eta$ each number $A_{\kappa}/Det\mathbf{A}$ must be an exact division or each $A_{\kappa}$ must be exactly dividable by $Det\mathbf{A}$.
By Lemma 3.10 the adjoint matrix $\mathbf{A}^{*}$ of the matrix $\mathbf{A}$ can be determined, so that numbers $A_{\kappa}$ for $\kappa=1,2,\ldots,\eta$ can also be determined. Particularly we have
$A_{1}=\sum_{\zeta=1}^{s}2^{\eta_{\zeta-1}+\sigma(\zeta-1)}3^{\eta-\eta_{\zeta}}(3^{\eta_{\zeta}-\eta_{\zeta-1}}-2^{\eta_{\zeta}-\eta_{\zeta-1}})$
$\sum_{\zeta=1}^{s}2^{\eta_{\zeta}+\sigma(\zeta-1)}3^{\eta-\eta_{\zeta}}(2^{\sigma(\zeta,\zeta)}-1)+Det\mathbf{A}.$
By Lemma 3.11 or by Lemma 3.12 $A_{1}$ can not be exactly dividable by $Det\mathbf{A}$ when $0<s<\infty$. Thus there exits no positive odd integer solution of the matrix equation $\mathbf{A}\mathbf{X}=\mathbf{B}$. It implies that the assumption $r_{c}(m,t,i,s)=1$ does not hold. Then there must be $r_{c}(m,t,i,s)\neq 1$ when $0<s<\infty$.
Hence no nontrivial Collatz cycle exists in $S_{c}(s)$ where $0<s<\infty$.
@@
The adjoint matrix $\mathbf{A}^{*}$ of a $\eta\times\eta$ nonsingular matrix $\mathbf{A}$ defined by Definition 1.9 can be determined where $Det\mathbf{A}=3^{\eta}-2^{\eta+\sigma(s)}$, $\eta=\eta_{s}$, $\eta_{0}=0$,
$\sigma(\zeta)=\sum_{j=1}^{\zeta}r_{j}\texttt{, }\sigma(\xi,\zeta)=\sum_{j=\xi}^{\zeta}r_{j}\texttt{, and }\eta_{\xi}=\sum_{j=0}^{\xi-1}(m_{j}-i_{j}+1)\texttt{ where }0\leq i_{j}\leq m_{j}$
for $\xi=1,2,\cdots,s$.
Let $A_{\kappa}=\sum_{j=1}^{\eta}A_{j,\kappa}$ for $\kappa=1,2,\ldots,\eta$. Then there are
(3.5) $A_{\kappa}=2^{\eta-\kappa+1+\sigma(s)}3^{\kappa-1}[1-(2/3)^{\kappa-1}]+3^{\eta}[1-(2/3)^{\eta_{1}-\kappa+1}]$
$\sum_{\zeta=1}^{s-1}2^{\eta_{\zeta}+1-\kappa+\sigma(\zeta)}3^{\eta-\eta_{\zeta}+\kappa-1}[1-(2/3)^{\eta_{\zeta+1}-\eta_{\zeta}}]\texttt{ for }\kappa=1,2,\ldots,\eta_{1}$
and for $\xi=1,2,\ldots,s-1$ and $\kappa=\eta_{\xi}+1,\eta_{\xi}+2,\ldots,\min(\eta_{\xi+1},\eta)$
(3.6) $A_{\kappa}=\sum_{\zeta=0}^{\xi}2^{\eta-\kappa+\eta_{\zeta}+1+\sigma(\xi+1,s)+\sigma(\zeta)}3^{\kappa-\eta_{\zeta}-1}[1-(2/3)^{\min(\eta_{\zeta+1},\kappa-1)-\eta_{\zeta}}]$
$3^{\eta}[1-(2/3)^{\eta_{\xi+1}-\kappa+1}]+\sum_{\zeta=\xi+1}^{s-1}2^{\eta_{\zeta}+1-\kappa+\sigma(\xi+1,\zeta)}3^{\eta-\eta_{\zeta}+\kappa-1}[1-(2/3)^{\eta_{\zeta+1}-\eta_{\zeta}}].$
Particularly we have
(3.7) $A_{1}=\sum_{\zeta=1}^{s}2^{\eta_{\zeta-1}+\sigma(\zeta-1)}3^{\eta-\eta_{\zeta}}(3^{\eta_{\zeta}-\eta_{\zeta-1}}-2^{\eta_{\zeta}-\eta_{\zeta-1}})$
$\sum_{\zeta=1}^{s}2^{\eta_{\zeta}+\sigma(\zeta-1)}3^{\eta-\eta_{\zeta}}(2^{\sigma(\zeta,\zeta)}-1)+Det\mathbf{A}.$